1. This 2 kg hammer is 0.4 m up. What is it's PE?

2. What is the KE of a 1500 kg car going at suburban speed of 14m/s (about 50 km/h or 30 mph)?

3. The same car is now going at highway speed of 28m/s (About 100 km/h or 60 mph)?

4. A 1 kg meteorite strikes the Moon at 11km/s. How much KE is that?


1
Expert's answer
2021-05-27T01:10:02-0400

1. PE = m × g × h = 2 kg × 9.8 m/s2 × 0.4 m = 7.84 J

2. KE = 1/2 × m × V² = 1/2 × 1500 kg × (14 m/s)² = 147000 J = 147 kJ

3. KE = 1/2 × 1500 kg × (28 m/s)² = 588000 J = 588 kJ

4. KE = 1/2 × m × V² = 1/2 × 1 kg × (11 km/s)² = 1/2 × 1 kg × (11000 m/s)² = 60500000 J = 60.5 MJ